Find the solutions to the following systems of equations:
(a){x^2+y^2= 100
y−x= 2
(b){y−log base9(x+ 1) = 0
y−log base9(x) = 1/2.

Found 2 solutions by rapture, stanbon:
Answer by rapture(86)

Post one question at a time for a faster reply.

I will partially answer question 1.

x^2 + y^2 = 100...Equation A
y = x + 2...Equation B

x^2 + (x + 2)^2 = 100...Solve this equation for x.

Then plug each value of x that you found into Equation B to find your y-values.

Understand?


Answer by stanbon(75887)

Find the solutions to the following systems of equations:
(a){x^2+y^2= 100
y−x= 2
----
x^2 + (x+2)^2 = 100
x^2 + x^2 + 4x -96 = 0
x^2 + 2x - 48 = 0
x = 6 or x = -8
If x = 6, y = -9
If x = -8, y =
---------------------------
(b){y−log base9(x+ 1) = 0
y−log base9(x) = 1/2.
-----
Substitute for "y":
log9(x)+(1/2) - log9(x+1) = 0
-----
log9[x/(x+1)] = -1/2
x/(x+1) = 9*(-1/2)
x/(x+1) = 1/3
-----
3x = x+1
2x = 1
x = 1/2
--------------
Cheers,
Stan H.
-------------